# Enhanced Robot Arm at the Edge with NLP and Vision Systems

### v1

@@ IV Conclusion and Future Work
This study has presented a ”proof of concept” for an assistive robot arm system enhanced with edge computing, NLP, and computer vision capabilities. Our results demonstrate the feasibility and effectiveness of using LLMs with vision systems to interpret and execute complex commands through natural language processing. The integration of these technologies has shown promise in improving the intuitiveness and accessibility of assistive robotic systems, making them versatile to the nuanced needs of users with disabilities. Our experimental findings highlight the significant potential of combining LLMs with computer vision to accurately interpret user intents and manipulate objects based on verbal commands. This synergy between language understanding and visual perception paves the way for more responsive and versatile assistive robots that can perform a wider array of tasks with greater precision and reliability.
Despite the successes, this study has identified several areas for future development. One of the limitations identified is the lack of feedback control in our current system. Future work will focus on integrating feedback mechanisms to allow the robot to adjust its actions in real-time based on environmental feedback, enhancing task execution accuracy and reliability. Building on the current proof of concept, future iterations include implementing our system on a more advanced Kinova 6 DOF assistive robot arm. This will allow us to explore the capabilities of our system in handling more complex tasks and environments. To ensure our system meets the needs of its intended users, comprehensive user studies will be conducted. These studies will involve individuals with disabilities to gather direct feedback on the system’s performance, usability, and areas for improvement. The current study has revealed certain limitations in the LLM’s ability to interpret complex or context-heavy commands. Ongoing research will be dedicated to refining the LLM’s understanding and processing capabilities, possibly through the development of a dedicated, localized LLM tailored specifically for robotic control applications. To enhance the system’s reliability and accessibility, further development will aim at achieving full network independence. By localizing all processing tasks, we aim to ensure functionality in environments with limited or no internet access, increasing versatility. Upgrading the vision pipeline, potentially to more advanced frameworks like YOLO, will be explored to overcome current limitations in object detection and interaction.

### v2

@@ IV Conclusion and Future Work
This study has presented a proof of concept for an assistive robot arm system enhanced with edge computing, NLP, and computer vision capabilities. Our results demonstrate the feasibility and effectiveness of using LLMs with vision systems to interpret and execute complex commands through natural language processing. The integration of these technologies has shown promise in improving the intuitiveness and accessibility of assistive robotic systems, making them versatile to the nuanced needs of users with disabilities. This synergy between language understanding and visual perception paves the way for more responsive and versatile assistive robots that can perform a wider array of tasks with greater precision and reliability.
Despite the successes, this study has identified several areas for future development. One of the limitations identified is the lack of feedback control in our current system. Future work will focus on integrating feedback mechanisms to allow the robot to adjust its actions in real-time based on environmental feedback, enhancing task execution accuracy and reliability. Building on the current proof of concept, future iterations include implementing our system on a more advanced Kinova 6 DOF assistive robot arm. This will allow us to explore the capabilities of our system in handling more complex tasks and environments. To ensure our system meets the needs of its intended users, comprehensive user studies will be conducted. These studies will involve individuals with disabilities to gather direct feedback on the system’s performance, usability, and areas for improvement. The current study has revealed certain limitations in the LLM’s ability to interpret complex or context-heavy commands. Ongoing research will be dedicated to refining the LLM’s understanding and processing capabilities, possibly through the development of a dedicated, localized LLM tailored specifically for robotic control applications. To enhance the system’s reliability and accessibility, further development will aim to achieve full network independence. By localizing all processing tasks, we aim to ensure functionality in environments with limited or no internet access, increasing versatility. Upgrading the vision pipeline, potentially to more advanced frameworks like YOLO, will be explored to overcome current object detection and interaction limitations.
Acknowledgement
This work has been partially supported by NSF Award # 2201536.